• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Saha

    Log entry time 16:02:19 on August28,2009

    Entry number 286601

    keyword=Shift Summary

    Shift Summary

    0800: Move to LH2 target
    1 uA beam with raster ON
    LT1 ~ 5k, RT1 ~ 6 k. Prescales S1 = 1.
    Raster size is 3.5 x 4

    Move to BeO target to change raster size to 2.5 x 2.5
    Raster current is 3.6 x 3.2
    Harp Scans, with and without raster
    Beam size at 03B: x = 270 um, y = 140 um

    Move to Empty Target.
    Going to 100 uA
    Set both HRS to 3.1416 GeV/c
    Switch to Integrative mode. Switch VDC hv OFF.
    Do BCM and BPM calibrations Luis Mercado & Kent Paschke

    Move to Carbon Target.
    Carbon Calibration runs with 70 uA

    Move to LH2 Cryo Target
    Beam set to 20 uA
    Going to 40 uA

    Going to Carbon Target
    Beam at 20 uA

    Beam down since 3 pm